hep.ucsb.edu

download report

Transcript hep.ucsb.edu

Update of US Testing Status
Anthony Affolder
On behalf of the US testing group
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
1
Fine Tuning of Fault Finding Cuts
•Combined measurements of all
modules produced with old hybrids
(100 W resistor to inverter)
•Fit Gaussian to central core of each
distribution
Old Hybrids
Channels Affected
By Change in Cut
 Found 5 s region
 All cuts now used (but 1) outside
of this region
 Low noise requirement too tight
• 2.8 s Deconvolution Mode
• 3.2 s Peak Mode
•Led to 33 false one sensor open
flags
•Re-applied fault finding with 5 s
requirements
• Ni > 1.55 (Deconvolution)
• Ni > 1.2 (Peak)
Module Testing Meeting, Dec. 11, 2003
•Removes false flagging while
not missing any new real faults
•This tuning should be done for
each module type
• Root combination software
available from UCSB
US Testing Status-Anthony Affolder
2
Comparison of Modules With
Old Vs. Final Hybrids
•Resistor to inverter stage changed
from 100 W to 50 W
•Good News
Old Hybrids
 A small number of each type of
TOB modules made
 All types have similar
distributions of cut variables
• Gives hope to idea that only
two sets of cuts needed
(1 sensor modules and 2
sensor modules)
•Bad News
 The noise (or gain maybe)
appears to be 5-10% lower
 May have to re-tune again
 Hopefully, gain changed due to
resistor switch
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
Final Hybrids
3
Newest CMN Problem Module (1051)
 Well within old selection criteria
 No large addition increase in
current during module assembly
 Old sensors
• 30210320274206
• 30210320274214
30200020001051
3000
2500
Bias Current (nA)
• Last SS6 module built using one
sensor with 1.2 mA extra current
(450 nA vs 1700 nA) in UCSB reprobing at 450 V.
2000
Current(DB)
Current(probing)
1500
Current(Bias Only)
Current (Bonded)
1000
500
0
0
100
200
300
Voltage
400
500
• CMN seen in chip 46 with
extremely high noise in channels
423-424
 Sensor flaw seen between two
channels
• Not clear if flaw cause of
problem
 Begins at 400 V where database
and measured bias current
diverge
• ~0.5 mA difference
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
4
New CMN Problem Module (1051)
• Module tested at slightly
elevated voltage to measure
effect as function of current
 Bias current 3.7 mA, < 2 mA
more than expected from
database
• For first half of chip, CM
subtracted noise a factor of
~1.75 higher than typical noise.
 A very little amount of microdischarge can cause the CM
subtraction algorithm not to
work properly
• CM subtraction algorithm
used is same as LT, and
test beam software
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
5
IV Test Results (UCSB)
Probed Current @ UCSB (400 V) – QTC Measurement (400 V)
Sensors
>2
mA
>5
mA
>10
mA
>20
mA
>100
mA
< -2
mA
<-5
mA
<-10
mA
OB2 (’00-01)
15%
9%
8%
5%
1%
8%
3%
1%
OB1 (’00-01)
6%
3%
3%
3%
3%
3%
0%
0%
OB2 (’02)
OB2 (’03)
3%
0%
3%
0%
0%
0%
0%
0%
0%
0%
2%
0%
2%
0%
0%
0%
• Environmental conditions tightly controlled
 Temperature 23.1-23.8 C
 RH < 30% at all times
• An increase greater than 5 mA can cause CMN
• Much better results with newer OB2 sensors (2002)
• None of the 20 newest (2003) OB2 sensor show any increase in bias
current!!!
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
6
IV re-probing (FNAL)
•FNAL has begun extensive reprobing program
30
25
I Fermi (microA)
 See R. Demaria’s talk in
sensor meeting
 Plan on re-probing all
sensors received so far
Idb vs Ifermi @ 400V - 2002
10
0
0
 24% are Grade B (>1.5 mA)
 >15% have >1.5 mA increase
in bias current relative to
QTC measurements
• 81 sensors from 2003 reprobed so far
15
5
•328 sensors already re-probed
5
10
15
20
25
30
I Production DB (microA)
Idb vs Ifermi @ 400V - 2003
30
25
20
I Fermi (microA)
• Earlier indications are that
the agreement between
QTC and re-probing
improved in 2003 sensors
20
15
10
5
0
0
5
10
15
20
25
30
I Production DB (microA)
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
7
CMN vs Batch
35
30
25
Number of
Sensors
20
15
10
5
0
138 143 147 148 150 202 203 216 218 219 220 318 323
Production Week
•Sensors which cause CMN are fairly evenly distributed throughout
production years 2001-2002
•Early indications are that 2003 may be better
 Extremely low statistics
 Only low bias current sensors used
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
8
Study of Common Mode
• The common mode point is
calculated event-by-event for
groupings of 32 channels
 The spectra of the common
mode is fit for groupings within
a chip with CMN problems
• Excluding the grouping with
high noise channel
• Spectra is fit with two Gaussians
 Central core plus tail
 Fit parameters are:
•
•
•
Fraction of events in tail
Width of central core
Width of tail
• Study how parameters vary with
current
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
9
Fit Result of Common Mode Point
Fraction of Events in Tail
• Fraction of events is flat with bias
current (~strip current)
• Width of central core increases with
bias current (~strip current)
• Width of tail increases with bias
current (~strip current) and may
flattens out at some current
0.60
Fraction
0.50
0.40
1011
1010
1014
1015
0.30
0.20
0.10
0.00
0
5
10
15
20
25
30
Bias Current (nA)
Width of Central Core
Width of Tail
1.20
0.80
1011
1010
1014
1015
0.60
0.40
0.20
0.00
0
5
10
15
20
25
30
Width (ADC)
Width (ADC)
1.00
20.00
18.00
16.00
14.00
12.00
10.00
8.00
6.00
4.00
2.00
0.00
1011
1010
1014
1015
0
5
Bias Current (nA)
Module Testing Meeting, Dec. 11, 2003
10
15
20
25
30
Bias Current (nA)
US Testing Status-Anthony Affolder
10
Studies with “Final” Sensors
• 10 modules produced with 20 “final” production
sensors
 OB2 produced in weeks 18-23 of 2003
 Extremely high quality
• Bias currents between 1-2 mA
• Only 2 pinholes not indicated in sensor database
• No high noise channels
• Effects of thermal cycling/strain studied
 Modules fixed to cold box plates at thermal contacts by screws
• Modules thermal cycled for about a week
 1 mm shims added under 1 or 2 contact points in order to
stress silicon
• ~3 more than offset in rod attachment points
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
11
Effects of Strain on Modules
• Modules attached to cold plate
with 4 screws through thermal
contacts
• To test the effects of twisting
modules, 1 mm shims added
under thermal contacts
 Bias current measured with 1 or
2 shims for all 10 modules
2
1 mm shim
1
• No change in current seen
Current (nA)
3020020020506 Strain Study
2000
1800
1600
1400
1200
1000
800
600
400
200
0
Current(clamps)
Current (screws)
Current (1-1mm shim)
Current (2-1mm shim)
0
100
200
300
400
500
Voltage
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
12
Modules Thermal Cycles
• Modules thermal cycled on modified cold plates
with/without shims
 168 module-hours with a total of 36 thermal cycles without
shims
 ~1000 module-hours with a total of ~200 thermal cycles with 2
shims
• No change in bias currents or noise seen
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
13
Single Rod Test Stand at UCSB
•Complete set of electronic
ready to test single rods
 Test box provides dry, dark,
and electrical isolated
environment
 Connects to Rod burn-in
chiller for cooling
•First production rod tested in
stand (J. Lamb & P. Gartung)
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
14
Rod Testing Results
•Faults clearly seen in rod using
new LT
ARCS
 Only opens on rod so far
•Laser gain differences add
complication to data analysis
 Fixed noise cuts will NOT
work due to 50% variation in
laser gain
 Hopefully laser gain can be
measured by header height
•Similar work on optimization of
calculation of pulse height &
peak time variable needed as in
module LT
Rod LT
 See P. Gartung’s talk
•More statistics needed in order
to know how best to test rod
Module Testing Meeting, Dec. 11, 2003
US Testing Status-Anthony Affolder
15